What real number(s) must be excluded from the domain of the given functions below?
1. f(x) = x+1 over -5x-10 (x+1/-5x-10) _______
2. f(x)= √4-x (square root 4-x) __________
3. f(x)= 2-x over x²-5x+6 (2-x/x²-5x+6) _________

Answer by rothauserc(4718)

1) (x + 1) / (-5x - 10) = (x + 1) / -5(x + 2)
exclude x = -2
:
2) square root(4 - x)
exclude all x > 4
:
3) (2 - x) / (x^2 - 5x + 6) = (2 - x) / ((x - 2) * (x - 3))
exclude x = 2, 3
: